# TrailBlazer: Trajectory Control for Diffusion-Based Video Generation

### 31 Dec 2023 v1

@@ 3.2 Spatial Cross Attention Guidance
The spatial cross attention modules are implemented in the denoising UNet module of Rombach et al. (2022). This module finds the cross attention between the query representation $\mathbf{Q}_{s}\in\mathbb{R}^{N_{F}\times d_{h}\times d}$ obtained from the SD latent $\mathbf{z}_{t}$, and the representations $\mathbf{K}_{s},\mathbf{V}_{s}\in\mathbb{R}^{N_{F}\times|W|\times d}$ of the $|W|$ prompt words from the text model, where $d$ is the feature dimension of the keys and queries. Usually $|W|\equiv 77$ when the text embedding model is CLIP Radford et al. (2021). The cross attention map Hertz et al. (2022) is then defined as $\mathbf{A}_{s}=\text{Softmax}(\mathbf{Q}_{s}\mathbf{K}_{s}^{T}/\sqrt{d})\in\mathbb{R}^{N_{F}\times d_{h}\times|W|}$,  [^2] Note that this is a “batch” matrix multiplication (e.g., the method torch.bmm in PyTorch Paszke et al. (2019)) that is $\mathbf{C}=\mathbf{A}\mathbf{B}\in\mathbb{R}^{b\times m\times n}$, where $\mathbf{A}\in\mathbb{R}^{b\times m\times p}$, and $\mathbf{B}\in\mathbb{R}^{b\times p\times n}$. Similarly, the transpose operation is $\mathbf{A}^{\top}\in\mathbb{R}^{b\times p\times m}.$ where $d_{h}\equiv w\times h$, defined by the spatial resolution height and width at the specific layer. Note for simplicity we omit the batch size and the number of the attention heads Vaswani et al. (2017) in our definition.
@@
$\text{W}_{s}(x,y)=\left\{\begin{matrix}c_{w},\quad(x,y)\in\mathcal{B}^{\prime}\\ 1,\quad\text{otherwise},\end{matrix}\right.$
$\text{S}_{s}(x,y)=\left\{\begin{matrix}c_{s}\ g(x,y),\quad(x,y)\in\mathcal{B}\\ 0,\quad\text{otherwise},\end{matrix}\right.$
@@
Given the set of indices of subject word prompts $\mathcal{I}$ and trailing maps $\mathcal{T}$, for each cross-activation component at location $(x,y)$ in $\mathbf{A}_{s}$ is modified as follows,
@@
where $\odot$ denotes the Hadamard (element-wise) product that scales the $x,y$ element of the cross-attention map $\mathbf{A}_{s}$ by the corresponding weight in $\text{W}_{s}(\cdot)$. The overall result is that the attention in the cross-attention map for the particular prompt word as well as the trailing maps, is stronger in the user-specified bbox region.
@@ 3.3 Temporal Cross-Frame Attention Guidance
To capture the temporal correlation in the video clip during training, a prevalent approach involves reshaping the latent tensor. This involves shifting the spatial information to the first dimension, a technique employed in VideoLDM Blattmann et al. (2023). This reshaping is done before passing the hidden activation into the temporal layers, allowing the model to learn about the “correlation” of spatial components through the convolutional layers. As shown in Fig. 3, the temporal attention map is obtained by $\mathbf{A}_{m}=\text{Softmax}(\mathbf{Q}_{m}\mathbf{K}_{m}^{T}/\sqrt{d})\in\mathbb{R}^{d_{h}\times N_{F}\times N_{F}}$, where $d_{h}$ is the spatial dimensions of this tensor, and $\mathbf{Q}_{m}\in\mathbb{R}^{d_{h}\times N_{F}\times d}$, and $\mathbf{K}_{m}\in\mathbb{R}^{d_{h}\times N_{F}\times d}$, respectively.
What is different from the spatial counterpart is that now $\mathbf{A}_{m}$ learns about the relation between the all activation components across frames. For instance, $\mathbf{A}_{m}^{(x,y,i,j)}$ denotes the activation at location $(x,y)$ between frame $i$ and frame $j$. We denote such tensors as $\mathbf{A}_{m}^{(i,j)}(x,y)$ to keep the notation consistency. As seen in our visual investigation (Fig. 2, last row), the background attention is higher when the cross frame attention compares the frames that are temporally far from each other, and the foreground attention is higher when the frames are closer in the video sequence.
@@
Here the normalized video temporal distance $d$ determines the level of the weight injection as the triangular window in time. Values $d\approx 0$ increase the activation inside the bbox. In contrast, when $d\approx 1$, the activation inside the box is reduced, thereby approximating the temporal “anti-correlation” effect seen in Fig. 2. The editing by $\text{S}_{m}(\cdot)$ is performed during the initial $N_{M}$ frames of the denoising process. See Fig. 9 and section 4.3 for an ablation on this parameter.
Then, similarly to Eq. 1, the temporal cross-frame attention map editing is performed as
$\mathbf{A}_{m}^{(i,j)}(x,y)\coloneqq\mathbf{A}_{m}^{(i,j)}(x,y)\odot\text{W}_{m}(x,y)+\text{S}_{m}(x,y)$ (2)
@@ 3.4 Scene compositing
The problem space becomes more complicated for video synthesis with more than one moving subject. Although the parameters $c_{s},c_{w}$ in Eq. 1 are specific to a particular subject, they indirectly affect the entire scene through the global denoising. Thus, the choice of these parameters for different subjects might interact and require a parameter search exponential in the number of subjects to find the best synthesis. If the prompt $\mathcal{P}$ and bbox $\mathcal{B}$ are in conflict then the result might be poor. For instance, a user may specify the moving of $\mathcal{B}$ from left to right associated with the prompt word “dog”, while the $\mathcal{P}$ is given as “a dog is sitting on the road”.
Considering the reasons above, we follow work such as Ma et al. (2023); Bar-Tal et al. (2023) that combine multiple subjects, each with their own prompt, during the latent denoising. These individual subjects are then composited into an overall image under the control of a global “composed” prompt. Our approach is illustrated in Fig. 4.
Figure 4: Scene Compositing. Given the set of latents generated from our system using a single bbox denoted as $\mathbf{z}^{(\text{ball})}_{t}$ and $\mathbf{z}^{(\text{dog})}_{t}$ for the case of prompts related to ball and dog. Then, the scene compositor produces a synthesis of multiple subjects with the complete prompt and the single subject latents. We refer reader to our supplementary video to view the implemented speed control of the dog.
Given a sequence of subject latents $\mathbf{z}^{(r)}_{t}$, each generated with different prompts using the core approach in Sec. 3.2 and Sec. 3.3, the final image is produced by a second denoising process in which the $N_{R}$ individual latents for the denoising step are composited using Eq. 3 for $N_{C}$ steps, followed by a global denoising in which the composited latent $\mathbf{z}_{t}$ is denoised using the complete (“composed”) prompt (pink box in Fig. 4),
$\mathbf{z}_{t}(x,y)\coloneqq\frac{1}{R}\sum_{r=0}^{N_{R}}w\,\mathbf{z}_{t}(x,y)+(1{-}w)\,\mathbf{z}^{(r)}_{t}\!(x,y),\quad\quad\forall t\in\{T,...,T{-}N_{C}\},\quad(x,y)\in\mathcal{B}_{r},$ (3)
where $w\in[0,1]$ determines the weight of linear interpolation between the specific subject latent $\mathbf{z}^{(r)}_{t}$ and the composed latent $\mathbf{z}_{t}$. It is formulated by considering the ratio of the current denoising timestep between $N_{C}$ and $T$,
$w=1-\frac{\big{(}N_{C}-(T-t)\big{)}}{N_{C}}.$
At the very beginning of the denoising step ($t=T$), the compositing fully prioritizes the subject latent $\mathbf{z}^{(r)}_{t}$ in each local region in the associated bbox $\mathcal{B}_{r}$. As $t$ decreases, $w$ gradually increases, giving higher priority to composed latent $\mathbf{z}_{t}$. This process concludes when $t=T-N_{C}$, resulting in $w=1$ which stops using the subject latent in the remaining denoising steps.
@@ 4 Experiments
Here we briefly present some experiments, evaluation, and limitations of our work. For full experiments and the implementation details we refer the reader to our supplementary and the project video for further examination. The figures here show an evenly spaced sampling of frames from the videos. Please zoom in to see the finer details of the results.
@@ 4.1 Main result
Fig. 5 shows our main result on trajectory control of a single subject. As a baseline, the right side of the figure shows Text2Video-Zero Khachatryan et al. (2023) [^3] We used the Huggingface implementation: https://huggingface.co/docs/diffusers/api/pipelines/text_to_video_zero. using the same prompts, and we used it without conditioning guidance (e.g., edge or depth maps) to provide a fair comparison. Text2Video-Zero accepts motion guidance in the form of a (x,y) translation vector. We set this vector to (8,0) to produce horizontal motion.
In Fig. 5, the left two sets are our results generated with linear interpolated bboxes starting at one side of the image and moving to the other boundary. It is evident that our result shows anatomically plausible motion of the subject. For instance, all subjects (e.g., cat, bee, astronaut, and clown fish) face in the direction that they move. This also applies for the other experiments in this paper. However, this does not generally happen in Text2Video-Zero as seen on the right. This is because they apply the the warping operation on the latent space directly, which simply translates the subject without re-orienting it. In addition it can be seen that background details are not preserved across frames.
@@
Figure 5: Main result: Rigid moving bbox. Transitioning the bbox movement from left to right is observed in the left set of four columns, while the transition from right to left occurs in the middle set of four columns. The set on the right is produced from Text2Video-Zero. The same prompt is used across each row. The video synthesis for each set of columns utilizes the same random seed and is generated from the identical prompt. (1st row): a cat walking on the grass field. (2nd row): A macro video of a bee pollinating a flower. (3rd row): An astronaut walking on the moon. (4th row): A clown fish swimming in a coral reef. The bold text represents the directed object.
@@
Fig. 6 illustrates dynamically changing the bbox size, producing an effect of the subject moving toward or away from the virtual camera. Similar to the Fig. 5, the bbox setup is from the top-left corner to the bottom-right corner. The dynamically changing bbox size is annotated with a green box. Note that the generated subjects share a desirable characteristic, that the subject naturally faces toward the virtual camera when the bbox transitions from small to large as seen in the sequences on the left, and vice versa on the right. The results also show a desirable perspective effect. Increasing or reducing the bbox size over time enables the synthesized object to produce the motion of “coming to” and “going away from” the camera. The red car also mimics motion blur and rising dust, simulating the effect of fast driving. We believe these effects arise naturally as a result of manipulating a model that was trained on video sequences rather than images. For example, while a video where a side-facing car appears bigger in the image is conceivable, for example if the camera is approaching the car from the side, it is much less common than videos where the car appears bigger due to it approaching.
Figure 6: Main result: Dynamic moving bbox. (Top row): A cat [X] on the road, where [X] denotes [walking $\rightarrow$ sitting] and [sitting $\rightarrow$ walking] for the left and right sequence, respectively. The bbox is initially moving, or keeping still in the left and the right sequence. (Middle row): A red car driving on the road. (Bottom row): The tiger walking on the street. The bbox used in middle and the bottom row are linearly interpolated with varied sizes. Please refer to the main text for more detail.
@@
Multi-subject synthesis is generally challenging, particularly when the number of objects exceeds two. We will delve into these limitations in Sec. 4.4. In Fig. 7, we present experiments with two subjects, a cat and a dog, guided by the green bbox in the sub-figure. The synthesis of the dog and cat in isolation is depicted in the top row, serving as a quality sanity check. Starting from the second row, we show the subsequent eight results combining environment prompts (e.g., “… on the moon”) after subject prompts (e.g., “A white cat and a yellow dog running…”). Each experiment demonstrates the flexibility of our method to synthesize subjects under varied environmental conditions. Notably, the interactions between the background and subjects appear plausible, as seen in reflections and splashes in the swimming pool case and consistent shadows across all samples. The results also show some artifacts such as extra limbs that are inherited from the underlying model. Indeed such artifacts are common in some diffusion models.
@@
Figure 7: Main result: Subjects compositing. Each set of the three sub-figures representing the first, middle, and the end frame of the synthesized video. The first row shows the video synthesis of the two subjects: “cat” and the “dog” guided by the bbox directed by the annotated arrows, respectively. Starting from the second row, each set of results show the varied post-fixed prompt.
@@ 4.2 Quantitative evaluation
Following the methodology in Blattmann et al. (2023); Hu and Xu (2023), we report Frechet Inception Distance Heusel et al. (2017) (FID), Inception Score (IS), and Kernel Inception Distance (KID) metrics against the CIFAR10 dataset Krizhevsky et al. on all images of video sequences, using the PyTorch implementation Obukhov et al. (2020). For a fair quantitative evaluation, we generated baseline results using Text2Video-Zero Khachatryan et al. (2023) without additional conditioning input. We set the motion vectors to $x=8,y=0$ and create a 24-frame video sequence as our baseline comparison. Subsequently, we manually annotated the beginning and end frames of their result with bboxes around the subject, and use these bboxes as keyframes to drive our system for comparison. We randomly sample 10 videos from each of the four prompts in Fig. 5, resulting in 960 testing images for each method. The result is summarized in Table. 1. As observed, our performance is roughly equivalent in terms of IS and KID, while our FID is significantly lower than that of Text2Video-Zero.
Method FID($\downarrow$) IS($\uparrow$) KID($\downarrow$)
Text2Video-Zero Khachatryan et al. (2023) 348.87 3.68 $\pm$ 0.19 0.19 $\pm$ 0.01
TrailBlazer 317.35 3.63 $\pm$ 0.25 0.18 $\pm$ 0.01
Table 1: Quantitative results.
Unfortunately there is no standardized quantitative measure for our task. For example, while the image samples from Text2Video-Zero in Fig. 5 appear visually reasonable, its synthesized video tends to exhibit temporal flickering, and the motion kinematics are not realistic (e.g., the cat moves to the left in Fig. 5 while consistently facing the observer). For a full and accurate assessment it is essential to see the video results. Please refer to our supplementary material and the accompanying video.
4.3 Ablations
We conduct ablation experiments on the number of trailing attention maps, and the number of temporal steps.
Trailing attention maps. Fig. 8 shows an ablation varying the number of trailing attention maps used in our spatial cross attention process, where the top row shows our method without trailing attention maps (e.g, $|\mathcal{T}|=0$) to $|\mathcal{T}|=30$. The guided bbox is identical to our earlier experiments in Fig. 5. It is observed that the astronaut remains static at the image center without the trailing attention maps. In contrast, the synthesis with a large number of trailing attentions can lead to failed results such as a flag rather than the intended astronaut. A good number of edited trailing attention maps is between $|\mathcal{T}|=10$ and $|\mathcal{T}|=20$.
Temporal attention editing. We further show an ablation test in Fig. 9 with varied number of temporal attention editing steps. We take the case of the astronaut from Fig. 8 with $|\mathcal{T}|=10$, and set $N_{M}=0$ (no editing steps), and $N_{M}=10$. The result with $N_{M}=0$ shows a red blob moving from left to right. The value $N_{M}=10$ gives satisfactory result on the astronaut, but the background along the bbox path is missing. From these results we see that a reasonable balance between spatial and the temporal attention editing must be maintained, while extreme values of either produce poor results. An intermediate value such as $N_{M}=5$ (used in most of our experiments) produces the desired result of an astronaut moving over a moon background.
4.4 Limitations
Our method shares and inherits common failure cases of the underlying diffusion model. Notably, at the time of writing, models based on CLIP and Stable Diffusion sometimes generate deformed objects and struggle to generate multiple objects and correctly assign attributes (e.g. color) to objects. We show some failures in Fig. 10. For instance, we requested a red jeep driving on the road but the synthesis shows it sinking into a mud road. The panda example shows the camera moving instead of the panda itself. The red car has implausible deformation, and Darth Vader’s light saber turns into a surf board. The length of the resulting video clips is restricted to that produced by the pre-trained model (24 images in the case of ZeroScope). This is not a crucial limitation, as movies are commonly (with some exceptions!) composed of short “shots” of several seconds each. The bounding box guides object placement without precisely constraining it. This is an advantage as well, however, since otherwise the user would have to specify the correct x-y aspect ratio for objects, a complicated task for non-artists.
Figure 10: Failure cases. Prompts used from (a) to (d): “A red jeep driving on the road”, “A red car driving on the highway”, “a panda eating bamboo”, and “Darth Vader surfing in waves”, where the bold prompt word is the directed subject.

### 08 Apr 2024 v2

@@ 3.2 Spatial Cross Attention Guidance
The spatial cross attention modules are implemented in the denoising UNet module of Rombach et al. (2022). This module finds the cross attention between the query representation $\mathbf{Q}_{s}\in\mathbb{R}^{N_{F}\times d_{h}\times d}$ obtained from the SD latent $\mathbf{z}_{t}$, and the representations $\mathbf{K}_{s},\mathbf{V}_{s}\in\mathbb{R}^{N_{F}\times|W|\times d}$ of the $|W|$ prompt words from the text model, where $d$ is the feature dimension of the keys and queries. Usually $|W|\equiv 77$ when the text embedding model is CLIP Radford et al. (2021). The cross attention map Hertz et al. (2022) is then defined as $\mathbf{A}_{s}=\text{Softmax}(\mathbf{Q}_{s}\mathbf{K}_{s}^{T}/\sqrt{d})\in\mathbb{R}^{N_{F}\times d_{h}\times|W|}$, [^2] Note that this is a “batch” matrix multiplication (e.g., the method torch.bmm in PyTorch Paszke et al. (2019)), that is $\mathbf{C}=\mathbf{A}\mathbf{B}\in\mathbb{R}^{b\times m\times n}$, where $\mathbf{A}\in\mathbb{R}^{b\times m\times p}$, and $\mathbf{B}\in\mathbb{R}^{b\times p\times n}$. Similarly, the transpose operation is $\mathbf{A}^{\top}\in\mathbb{R}^{b\times p\times m}.$ where $d_{h}\equiv w\times h$, defined by the spatial resolution height and width at the specific layer. For simplicity we omit the batch size and the number of attention heads Vaswani et al. (2017) in our definition.
@@
$\text{S}_{s}(x,y)=\left\{\begin{matrix}c_{s}\ g(x,y),\quad(x,y)\in\mathcal{B}\\ 0,\quad\text{otherwise},\end{matrix}\right.\qquad\text{W}_{s}(x,y)=\left\{\begin{matrix}c_{w},\quad(x,y)\in\mathcal{B}^{\prime}\\ 1,\quad\text{otherwise},\end{matrix}\right.$ (1)
@@
Given the set of indices of subject prompt words $\mathcal{I}$ and trailing maps $\mathcal{T}$, each cross-activation component at location $(x,y)$ in $\mathbf{A}_{s}$ is modified as follows:
@@
where $\odot$ denotes the Hadamard (element-wise) product that scales the $x,y$ element of the cross-attention map $\mathbf{A}_{s}$ by the corresponding weight in $\text{W}_{s}(\cdot)$. The result is that the attention in the cross-attention map for the particular prompt word as well as the trailing maps, is stronger in the user-specified bbox region.
@@ 3.3 Temporal Cross-Frame Attention Guidance
To capture the temporal correlation in the video clip during training, a prevalent approach involves reshaping the latent tensor. This involves shifting the spatial information to the first dimension, a technique employed in VideoLDM Blattmann et al. (2023). The reshaping is done before passing the hidden activation into the temporal layers, allowing the model to learn about the “correlation” of spatial components through the convolutional layers. As shown in blue section in Fig. 3, the temporal attention map is obtained by $\mathbf{A}_{m}=\text{Softmax}(\mathbf{Q}_{m}\mathbf{K}_{m}^{T}/\sqrt{d})\in\mathbb{R}^{d_{h}\times N_{F}\times N_{F}}$, where $d_{h}$ is the spatial dimensions of this tensor, $\mathbf{Q}_{m}\in\mathbb{R}^{d_{h}\times N_{F}\times d}$, and $\mathbf{K}_{m}\in\mathbb{R}^{d_{h}\times N_{F}\times d}$.
What is different from the spatial counterpart is that now $\mathbf{A}_{m}$ learns about the relation between the correlated components across all frames. For instance, $\mathbf{A}_{m}^{(x,y,i,j)}$ denotes the correlation at location $(x,y)$ between frame $i$ and frame $j$. We denote such tensors as $\mathbf{A}_{m}^{(i,j)}(x,y)$ to keep the notation consistent. As seen in our visual investigation (Fig. 2, Right), the background attention is higher when the cross frame attention compares the frames that are temporally far from each other, and the foreground attention is higher when the frames are temporally closer in the video sequence.
@@
Here the normalized video temporal distance $d$ determines the level of the weight injection as a triangular window in time. Values $d\approx 0$ increase the activation inside the bbox. In contrast, when $d\approx 1$, the activation inside the box is reduced, approximating the temporal “anti-correlation” effect seen in Fig. 2. The editing by $\text{S}_{m}(\cdot)$ is performed during the initial $N_{M}$ steps of the denoising process.
Then, similarly to Eq. 2, the temporal cross-frame attention map editing is,
$\mathbf{A}_{m}^{(i,j)}(x,y)\coloneqq\mathbf{A}_{m}^{(i,j)}(x,y)\odot\text{W}_{m}(x,y)+\text{S}_{m}(x,y),$ (3)
@@ 3.4 Scene compositing
The problem space becomes more complicated for video synthesis with more than one moving subject. Although the parameters $c_{s},c_{w}$ in Eq. 2 are specific to a particular subject, they indirectly affect the entire scene through the global denoising. Thus, the choice of these parameters for different subjects might interact and require a parameter search in the number of subjects to find the best synthesis. If the prompt $\mathcal{P}$ and bbox $\mathcal{B}$ are in conflict then the result might be poor. For instance, a user may specify motion of $\mathcal{B}$ from left to right associated with the prompt word “dog”, while $\mathcal{P}$ is given as “a dog is sitting on the road”. In fact, the dog moves in accordance with the configured bbox, either walking or running.
Considering the reasons above, we follow work such as Ma et al. (2023); Bar-Tal et al. (2023) that combine multiple subjects, each with their own prompt, during the latent denoising. The latents $\mathbf{z}^{(r)}_{t}$ for the $r$-th subject are then composited into an overall image latent $\mathbf{z}_{t}$ under the control of a “composed” prompt, as illustrated in Fig. 4 and formulated as,
Figure 4: Scene Compositing. Given the set of latents generated from our system using a single bbox denoted as $\mathbf{z}^{(\text{ball})}_{t}$ and $\mathbf{z}^{(\text{dog})}_{t}$ for the case of prompts related to ball and dog. Then, the scene compositor (SC) produces a synthesis of multiple subjects with the complete prompt and the single subject latents. We refer reader to our supplementary video to view the implemented speed control of the dog.
$\mathbf{z}_{t}(x,y)\coloneqq\frac{1}{R}\sum_{r=0}^{N_{R}}w\,\mathbf{z}_{t}(x,y)+(1{-}w)\,\mathbf{z}^{(r)}_{t}\!(x,y),$ (4)
where $\forall t\in\{T,...,T{-}N_{C}\},\quad(x,y)\in\mathcal{B}_{r}$, where $w\in[0,1]$ determines the weight of linear interpolation between the specific subject latent $\mathbf{z}^{(r)}_{t}$ and the composed latent $\mathbf{z}_{t}$. It is formulated by considering the ratio of the current denoising timestep between $N_{C}$ and $T$, such that $w=1-\big{(}N_{C}-(T-t)\big{)}/N_{C}$. At the beginning of the denoising process (so at $t=T$), the compositing fully prioritizes the subject latent $\mathbf{z}^{(r)}_{t}$ in each local region in the associated bbox $\mathcal{B}_{r}$. As $t$ decreases, $w$ gradually increases, giving higher priority to composed latent $\mathbf{z}_{t}$. This process concludes when $t=T-N_{C}$, resulting in $w=1$, which stops using the subject latent in the remaining denoising steps.
@@ 4 Experiments
Here we briefly present some experiments and quantitative evaluations of our work. Please see our supplementary materials and the project video for full experiments, including implementation details, limitations, ablations, comparisons, and finer details. The figures show an evenly spaced temporal sampling of frames from the videos.
@@ 4.1 Main result
Fig. 5 shows our main result on trajectory control of a single subject. We compare TrailBlazer to T2V-Zero Khachatryan et al. (2023) and Peekaboo Jain et al. (2023) using the same prompts, without conditioning guidance (e.g., edge or depth maps) to provide a fair comparison. T2V-Zero accepts motion guidance in the form of an (x,y) translation vector. We set this vector to (8,0) to produce horizontal motion. More detailed visual comparisons with Peekaboo under extreme conditions are depicted in Sec. 9 of the supplementary materials. The presented results for each method are visually selected as the best of a pool of 10 experiments conducted with different random seeds.
In Fig. 5, the results are generated from linearly interpolated bboxes starting at the left of the image and moving to the right. The results from TrailBlazer demonstrate anatomically plausible motion of the subject and a more accurate fitting of the subject within the bbox. Further, all subjects (e.g., cat, bee, astronaut, and clown fish) face in the direction that they move. However, this is not a common occurrence in T2V-Zero, as they directly apply the editing operation on the diffusion latents. This approach merely translates the subject without re-orienting it. Although the synthesized subject’s motion generally follows the bbox in Peekaboo, it does not fit the bbox well. Occasionally, artifacts may emerge, such as a rectangular object following the astronaut. Moreover, our synthesized background exhibits better visual quality. In the competing methods, the background often appears plain, blurry, or lacks detail behind the subject (e.g., the area behind the subject path in Peekaboo).
@@
Figure 5: Main result: Rigid bbox moving from left to right. TrailBlazer and Peekaboo use identical bboxes, while T2V-Zero uses the corresponding motion vector instead. The same prompt is used across each method. The four prompts used (clockwise from top left): An astronaut walking on the moon; A macro video of a bee pollinating a flower; A clown fish swimming in a coral reef; A cat walking on the grass field. The bold text represents the directed object.
@@
Fig. 6 illustrates the speed control, and dynamically changing the bbox size producing an effect of the subject moving toward or away from the virtual camera. In the top two rows of comparisons, the bbox setup is between the top-left corner and the bottom-right corner. The dynamically changing bbox size is annotated with a green box as illustrated on the left. Note that the generated subjects share the desirable characteristic that the subject naturally faces toward the virtual camera when the bbox transitions from small to large as seen in the top sequence and vice versa in the second sequence from the top. The results also show a desirable perspective effect. Increasing or reducing the bbox size over time causes the synthesized object to produce the motion of “coming to” and “going away from” the camera as shown in the tiger example. We believe these effects arise naturally as a result of manipulating a model that was trained on video sequences rather than images. Peekaboo’s tiger fails to produce these perspective effects when guided with identical bboxes.
Furthermore, our method adeptly manages fast motion, outperforming Peekaboo in this regard. This is evident in the third row of both subfigures of Fig. 6, which illustrate a cat rapidly running from one side to the other multiple times in the video clip, where $N_{f}=24$. More precisely, the bbox is initially positioned at the left (1st keyframe), after which it is moved to the right (2nd keyframe), then left (3rd keyframe), right (4th keyframe), and left (5th keyframe).
Figure 6: Main result: Dynamic moving bbox. (Top/Middle row): The tiger walking on the street. (Bottom row): The cat running on the grass. The first column illustrates the bbox keyframes in the squared layout, where the green bbox is guided by the almond-colored motion vector. Note that there are five keyframes in the third row of each subfigure, with the bbox located on the left at the initial keyframe. The bbox used in top and the middle row are linearly interpolated with varied sizes. The bbox used in the last row has a static size with 5 keyframes moving back-and-forth. Please refer to the main text for more detail.
@@
Multi-subject synthesis is generally challenging, particularly when the number of objects exceeds two. We delve into this issue in the supplementary materials. In Fig. 7, we present experiments with two subjects, a cat and a dog, guided by the green bbox in the sub-figure. The synthesis of the dog and cat in isolation is depicted in the top row on the left, serving as a sanity check with annotated image frame. We also show six results combining environment prompts (e.g., “… on the moon”) after composed prompt (e.g., “A white cat and a yellow dog running…”). Each experiment demonstrates the flexibility of TrailBlazer in synthesizing subjects under varied environmental conditions. Notably, the interactions between the background and subjects appear plausible, as seen in reflections and splashes in the swimming pool case and consistent shadows across all samples. The results also show some artifacts such as extra limbs that are inherited from the underlying model.
@@
Figure 7: Main result: Subjects compositing. Each set of the three sub-figures representing the first, middle, and the end frame of the synthesized video. The first row on the left with annotated frame shows the video synthesis of the two subjects: “cat” and the “dog” guided by the bbox directed by the annotated arrows, respectively. Next, each set of results show the varied post-fixed prompt.
@@ 4.2 Quantitative evaluation
Following the methodology in Blattmann et al. (2023); Hu and Xu (2023); Jain et al. (2023), we report Fréchet Inception Distance Heusel et al. (2017) (FID), Fréchet Video Distance (FVD), Inception Score (IS), Kernel Inception Distance (KID), mean intersection of union (mIoU), and CLIP similarities (CLIPSim) metrics against the random selected 400 videos in AnimalKingdom dataset Ng et al. (2022) on all images of video sequences. As described in the supplementary materials, we evaluate both methods using the prompt set published in Jain et al. (2023). The mIoU evaluation utilizes the OWL-ViT-large open-vocabulary object detector Minderer et al. (2022) to obtain the bbox of the synthesized subject.
For a fair quantitative evaluation, we generated baseline results using Peekaboo Jain et al. (2023) and T2V-Zero Khachatryan et al. (2023) without additional conditioning input. Both TrailBlazer and Peekaboo share the same keyframed bbox, the motion vectors are used for T2V-Zero depending on the tasks below, and we use a 24-frame video sequence as our baseline comparison. We conducted two experiments with the associated random keyframing for our work: Static bbox, and Dynamic bbox.
The bboxes in the Static bbox experiments are constant across all keyframes, where the top left corner is randomly generated in the second quadrant, and the width and height is randomly selected between 25% to 50% of the image resolution. This experiments mainly evaluate the method without considering the bbox motion. The result is summarized in Table. 1. As observed, our performance is roughly equivalent across all metrics, while our FVD is significantly lower than that of T2V-Zero and Peekaboo. Motion vector $x=0,y=0$ is used in T2V-Zero.
Method FID($\downarrow$) FVD($\downarrow$) IS($\uparrow$) KID($\downarrow$) mIoU($\uparrow$) CLIPSim($\uparrow$)
T2V-Zero Khachatryan et al. (2023) 198.45 2897.57 2.48 $\pm$ 0.31 4.39% - 31.54
Peekaboo Jain et al. (2023) 159.55 1521.27 2.27 $\pm$ 0.49 2.75% 0.23 31.31
TrailBlazer 182.28 1220.98 2.91 $\pm$ 0.54 3.51% 0.26 30.93
Table 1: Quantitative results for static bbox.
Table. 2 presents the results of the Dynamic bbox experiments. The bboxes were generated by randomly specifying 2 to 6 keyframes evenly in the video clip where the location is at the image boundary and its opposite as shown in Fig. 6. The height and width of the particular bbox is between 10% to 50%. Thus the location and the size of all bboxes are varied in a video clip. Motion vector $x=8,y=0$ is used in T2V-Zero.
In Table. 2, the notable improvement is our mIoU score compared to Peekaboo, which can be attributed to the capabilities demonstrated in Fig. 6, showcasing TrailBlazer’s proficiency in generating a perspective view with dynamically changing bboxes. In comparison to Peekaboo, TrailBlazer’s FID is better, while they exhibit better FVD. This discrepancy may be explained by the nature of the AnimalKingdom Ng et al. (2022) dataset, where creatures typically perform actions in a stationary setting (e.g., birds singing, animals walking). Notably, the running cat motion in Fig. 6 is generally absent in their dataset, contributing to the lower FVD score in our case. Our better FID score suggests that the individual frame quality in our video clip is better.
Method FID($\downarrow$) FVD($\downarrow$) IS($\uparrow$) KID($\downarrow$) mIoU($\uparrow$) CLIPSim($\uparrow$)
T2V-Zero Khachatryan et al. (2023) 153.36 2350.69 4.51 $\pm$ 1.36 3.01% - 31.32
Peekaboo Jain et al. (2023) 165.88 1571.04 2.16 $\pm$ 0.57 2.86% 0.25 31.92
TrailBlazer 148.18 1721.08 2.45 $\pm$ 0.61 3.12% 0.37 30.71
Table 2: Quantitative results for dynamic bbox.
In summary, the objective scores in Tables 1, 2 do not give a clear ordering of methods. However, recall that our goal is controlling movement. TrailBlazer achieves this, showing significantly better mIoU scores. Equally important, TrailBlazer shows improved subjective movement, with moving objects facing in plausible directions and having realistic motion. Please refer to our video.